OpenStudy (anonymous):

It takes Karina 3.5 hours to drive from Davenport to Hogsdale and then 2 hours and 10 minutes to drive from Hogsdale to Paqua. On the route that she took, it is 720 miles from Davenport to Paqua. What was her average driving speed for the trip?

mathslover (mathslover):

oh k so do you know what is average speed formula?

mathslover (mathslover):

@julialeigh I will not be able to help u unless u corporate me .. response to my quest.

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

no i dont know the formula

OpenStudy (anonymous):

@mathslover u there?

mathslover (mathslover):

3.5 hrs + 2 hrs + 10 minutes = 5 hours and 40 minutes = 5 hours and 2/3 hours Average speed = s / t =720 miles / (5 hours + 2/3 hours) =720 miles / (17/3 hours) = 127.06 miles/hours (approximation) http://wiki.answers.com/Q/If_Karina_takes_3.5_hrs_to_drive_from_davenport_to_hogsdale_and_then_2_hrs_and_10_mins_to_drive_from_hogsdale_to_paqua_It_is_720_miles_from_davenport_to_paqua_what_was_her_average_driving_speed

OpenStudy (anonymous):

Okay so how did you get 17/3 ? and the actual answer? Work shown would be appreciated I am trying to learn how to do the problems and show work.